ASYMMETRIC TOP ROTORS IN SUPERFLUID PARA-HYDROGEN NANO-CLUSTERS

\underline{TAO ZENG}, Department of Chemistry, University of Waterloo, Waterloo, ON, Canada, N2L 3G1; HUI LI, Institute of Theoretical Chemistry, State Key Laboratory of Theoretical and Computational Chemistry, Jilin University, 2519 Jiefang Road, Changchun 130023, People’s Republic of China; and PIERRE-NICHOLAS ROY, Department of Chemistry, University of Waterloo, Waterloo, ON, Canada, N2L 3G1.

We present the first simulation study of bosonic clusters doped with an asymmetric top molecule. A variation of the path-integral Monte Carlo method is developed to study a para-water (pH_{2}O) impurity in para-hydrogen (pH_{2}) clusters. The growth pattern of the doped clusters is similar in nature to that of the pure clusters. The pH_{2}O molecule appears to rotate freely in the cluster due to its large rotational constants and the lack of adiabatic following. The presence of pH_{2}O substantially quenches the superfluid response of pH_{2} with respect to the space fixed frame. We also study the behaviour of a sulphur dioxide (^{32}S^{16}O_{2}) dopant in the pH_{2} clusters. For such a heavy rotor, the adiabatic following of the pH_{2} molecules is established and the superfluid renormalization of the rotational constants is observed. The rotational structure of the SO_{2}-p(H_{2})_{N} clusters’ ro-vibrational spectra is predicted. The connection between the superfluid response respect to the external boundary rotation and the dopant rotation is discussed.